# Rigidity and nonexistence of CMC hypersurfaces in 5-manifolds

### v2

@@ 1.2. Main ideas
The strategy in this article follows from that in [chodosh-li-stryker] and [Hong24]. A common feature in the proofs of Theorem 1.8 and 1.12 is the estimate on volume growth of ends, using $\mu$-bubbles, a technique introduced by Gromov [gromovmububble]. The method in [chodosh-li-stryker] cannot extend to $5$-dimensions directly because one cannot control curvature terms from the second fundamental form, the ambient Ricci curvature and the positive Jacobi function in the second variation of the warped $\mu$-bubble constructed there. The novelty of our method is to introduce a parameter (the exponent of the warped function) which brings more freedom to exploit curvature conditions. The same strategy has been applied successfully in the study of CMC hypersurfaces in $\mathbb{H}^{4}$ [Hong24].
@@
In Theorem 1.9, consider $M^{4}\hookrightarrow(X^{5},g)$ a complete two-sided weakly stable minimal hypersurface.
@@
    To show the vanishing of $A_{M}$ and $\operatorname{Ric}_{X}(\eta,\eta)$, on the contrary we show that if $A_{M}+\operatorname{Ric}_{X}(\eta,\eta)$ is strictly positive somewhere, we can find a sequence of compactly supported test functions $\varphi_{i}$ such that $\int_{M}|\nabla\varphi_{i}|^{2}=o(1)$, thus contradicting the weak stability.
@@
    In the same spirit in [Cao-Shen-Zhu-infinitevolume], the fact that $M$ has at most one nonparabolic end follows from the Schoen–Yau inequality (just uses $\operatorname{biRic}_{X}\geq 0$) and the infinite volume of $M$ (guaranteed by bounded geometry). If we assume $\operatorname{Ric}^{X}_{3}\geq 0$, bounded geometry condition can be weaken to weakly bounded geometry condition; see [chodosh-li-stryker][Theorem 4.1] in detail.
@@
    As mentioned in [chodosh-li-stryker], it is difficult to prove the linear volume growth of the nonparabolic end. Similarly, we consider an exhaustion $\{\Omega_{l}\}_{l=1}^{\infty}$ of $E$ and use the theory of $\mu$-bubbles to show $\partial\Omega_{l}\cap E$ has controlled diameter. To control curvature terms in the second variation of the $\mu$-bubble, we introduce a parameter $k$ as the exponent of the warped function. Suitable choice of $k$ yields the desired diameter estimate; see Lemma 4.6 in detail. This is the step that needs the assumption of strictly positive $k\operatorname{-triRic}$.
@@
    We need to upgrade diameter control to volume control which needs a lower bound on the intrinsic Ricci curvature of $M$. It is achieved by combining $\operatorname{biRic}_{X}\geq 0$, curvature estiamtes in Lemma 2.4 and bounded geometry.
@@
    Finally, we need to construct balanced test function for weak stability in step (i). This is achievable due to the facts that we have at most one nonparabolic end and the nonparabolic end satisfies almost linear volume growth. In fact, we can prove the following general result which is of independent interest.
Proposition 1.13.
Let $M$ be a complete manifold with infinite volume and finite nonparabolic ends $\{E^{k}\}_{k=1}^{k_{0}}$. For a fixed scale $L$ and each $k$, let $E^{k}_{l}$ be the nonparabolic component of $E^{k}\setminus B_{(l+1)L}(p)$. Suppose that each $E^{k}$ satisfies almost linear volume growth, i.e., $E_{l}\cap B_{(l+1)L}(p)\leq C$ for a universal constant $C$. Then there exists a compact supported Lipschitz function $\varphi$ such that
$\int_{M}\varphi=0\ \text{and}\ \ -\int_{M}\varphi L(\varphi)<0$
for any operator $L=\Delta_{M}+q(x)$ where $q(x):M\rightarrow\mathbb{R}_{\geq 0}$ is not identically zero.
The proof of Theorem 5.2 follows the same strategy in [Hong24] by the first author. We sketch it as follows: The finite index assumption implies that any end $E$ of $M$ is stable outside a compact set. Combining the condition 5.1 with the stability inequality yields the exponential volume growth of the universal covering $\widetilde{E}$, which contradicts with the almost linear growth implied by Lemma 4.7. This is where condition (5.2) and universal subexponential end growth condition enter the proof.
We end the introduction with the following conjecture.
Conjecture: There is no complete noncompact constant mean curvature hypersurface $M$ in hyperbolic space $\mathbb{H}^{5}$ satisfying
$\frac{\sqrt{175}}{\sqrt{148}}>H_{M}>1\ \ \text{and}\ \ \operatorname{Ind}<\infty.$
@@ 1.3. Organization
In Section 2, we recall some preliminaries on curvature estimates and the ends of non-compact manifolds. In Section 3, we prove the one-end theorem under $\operatorname{biRic}_{X}\geq 0$ and bounded geometry assumption. We discuss the balanced warped $\mu$-bubble exhaustion result and the volume growth estimates in Section 4, and prove main results in Section 5. Appendix contains an example relevant to Theorem 1.9.
@@ 1.4. Acknowledgements
The first author is supported by a Science grant (YA24XKRC00010) by Beijing Jiaotong University. We would like to thank Prof. Xianzhe Dai, Prof. Martin Li, Prof. Guofang Wei and Prof. Rugang Ye for their interest in this work.
@@ 2.1. Curvature conditions
If $(X^{n+1},g)$ is a Riemannian manifold, we denote the Riemann curvature tensor, the sectional curvature, the Ricci curvature and the scalar curvature by $\operatorname{Rm}_{X}$, $\sec_{X}$, $\operatorname{Ric}_{X}$ and $R_{X}$ respectively. Let $M\hookrightarrow X^{n+1}$ be an immersion. We write $\operatorname{Ric}_{M}$ and $R_{M}$ to denote the Ricci curvature and scalar curvature of the pullback metric on $M$. We set the mean curvature to be the trace of the second fundamental form of the immersion.
@@
$\sec_{X}\leqslant\delta^{2},\quad\operatorname{inj}_{X}\geqslant\lambda$
@@
Let $(X^{n+1},g)$ be a Riemannian manifold and $\{e_{i}\}_{i=1}^{n+1}$ be an orthonormal basis for $T_{p}X$, $p\in X$. The bi-Ricci curvature is defined by
$\operatorname{biRic}_{X}(e_{1},e_{2}):=\sum_{i=2}^{n+1}\operatorname{Rm}_{X}(e_{1},e_{i},e_{i},e_{1})+\sum_{j=3}^{n+1}\operatorname{Rm}_{X}(e_{2},e_{j},e_{j},e_{2}).$
@@
Let $(X^{5},g)$ be a $5$-dimensional Riemannian manifold with bounded geometry and $\operatorname{biRic}_{X}\geq 0$. If $M^{4}\hookrightarrow(X^{5},g)$ is an immersed CMC hypersurface, then
$\operatorname{Ric}_{M}\geqslant-Cg,$
where the positive constant $C$ depends on $\sec_{X}$, $H_{M}$ and $|A_{M}|$.
@@
Note that bounded geometry and nonnegative bi-Ricci condition imply that
$|\sec_{X}|\leqslant 6\delta^{2}.$
Choose an orthonormal basis $\{e_{i}\}_{i=1}^{5}$ of $X^{5}$ at a point $p\in M$ with $e_{5}$ normal to $M$. The traced Gauss equation gives
$\operatorname{Ric}_{M}(e_{1},e_{1})=\sum_{i=2}^{4}\operatorname{Rm}^{X}_{i11i}+A_{M}(e_{1},e_{1})H_{M}-\sum_{j=1}^{4}A_{M}(e_{1},e_{j})^{2},$
@@
2.2. Finite index and curvature estimates
Recently, Chodosh, Li, Minter and Stryker proved that a two-sided complete stable minimal hypersurface $M^{4}\hookrightarrow\mathbb{R}^{5}$ is flat [Chodosh-Li-Minter-Stryker]. A standard blow-up argument yields curvature estimates for stable CMC hypersurfaces in non-compact $5$-manifolds.
@@ 2.2. Finite index and curvature estimates
Let $(X^{5},g)$ be a complete Riemannian manifold with bounded geometry and $\operatorname{biRic}\geqslant 0$. Then there is a constant $C$ such that every noncompact stable CMC hypersurface (two-sided if it is minimal) $M^{4}\hookrightarrow(X^{5},g)$ satisfies
$\sup_{q\in\Omega}|A_{M}(q)|\min\{1,d_{M}(q,\partial_{1}\Omega)\}\leqslant C$
for any compact subset $\Omega\subset M$ where $\partial_{1}\Omega=\partial\Omega\cap\mathring{M}.$
@@
Note that bounded geometry and $\operatorname{biRic}\geqslant 0$ implies the boundness of $\sec_{X}$. Proposition D.1 in [chodosh-li-stryker] shows that $Q$-weakly bounded geometry follows from $|\sec_{X}|\leqslant K$. Then, the proof is similar to that of [chodosh-li-stryker][Lemma 2.4] and [Hong24][Lemma 4.5] with minor modification, see also [Rosenberg-Souam-Toubiana]. As indicated in the proof of Lemma 4.5 in [Hong24], we only need to take care of the case when $M$ is not geodesically complete. ∎
2.3. Parabolicity and nonparabolicity
In this subsection, we recall some preliminary results [Hong24] on parabolicity and nonparabolicity of manifolds with noncompact boundary; see [chodosh-li-stryker] for detailed discussion on manifolds without boundary.
Let $M^{n}$ be a noncompact Riemannian manifold with boundary $\partial M$. Denote the outer unit normal vector of $\partial M$ by $\tau$. Choose a large compact subset $K$ containing $p\in\partial M$ such that the complement $M\backslash K$ has many disjoint components which are called ends. Boundary ends are those sharing boundary portion with $\partial M$ while interior ends only have boundary lying in $\partial K$. In dimension two, these two concepts are used to study the structure of capillary surfaces by the first author [hongcrelle].
Definition 2.5 ([Wu23]).
Denote a boundary end by $E$ and $\partial E$ consists of two parts: $\partial_{1}E:=\partial E\cap\partial K$ and $\partial_{0}E:=\partial E\cap\partial M$. $E$ is parabolic if there is no positive harmonic function $f\in C^{2,\alpha}(E)$, $10<\alpha<$ 1, so that
$f|_{\partial_{1}E}=1,\quad\partial_{\tau}f|_{\partial_{0}E}=0,\quad f|_{\mathring{E}}<1.$
Otherwise, we say that $E$ is nonparabolic and such $f$ is said to be a barrier function of $E$.
Remark 2.6.
The claimed regularity can be guaranteed by perturbing the intersection angle btween $\partial_{0}E$ and $\partial_{1}E$ to lie in $(0,\frac{\pi}{8})$ [Wu23]. In the sequel, we always assume that $\partial_{0}E$ meets $\partial_{1}E$ with a constant angle in $(0,\frac{\pi}{8})$.
We conclude this section by two useful characterizations of nonparabolicity.
Proposition 2.7 ([Wu23]Lemma 4.8).
A boundary end $E$ is nonparabolic if and only if the sequence of positive harmonic functions defined on $E_{i}:=E\cap B_{R_{i}}$, $R_{i}\to\infty$, satisfying
$f_{i}|_{\partial_{1}E}=1,\quad\partial_{\tau}f_{i}|_{\partial_{0}E}=0,\quad f_{i}|_{\partial B_{R_{i}}\cap E}=0$
converges uniformly and locally on compact subsets of $E$ to a barrier function $f$. Moreover, the barrier is minimal among all harmonic functions satisfying mixed boundary conditions, and has finite Dirichlet energy.
If the end $E$ is parabolic, the sequence constructed above converges to constant function $1$ and the Dirichlet energy goes to zero.
@@ 2.3. Parabolicity and nonparabolicity
A boundary end $\{E_{k}\}_{k=0}$ with length scale $L>0$ is nonparabolic if for every $k$, $E_{k}$ is an unbounded component of $M\backslash B_{kL+R_{0}}$ with nonparabolic boundary portion containing in $\partial M$ and $E_{k+1}\subset E_{k}$. Moreover, $E_{k}$ is nonparabolic.
An interior end is defined similarly.
3. One nonparabolic end for minimal hypersurface
In this section, following [Li-Wang-nonnegatively-curved-manifold], we prove that in a manifold $(X^{n+1},g)$ with $\operatorname{biRic}_{X}\geqslant 0$, the number of nonparabolic ends of a weakly stable minimal hypersurface $(M^{n},g)$ is at most one. This fact is known to experts. We include some details and adjustments for convenience of readers.

### v3

@@ 1.2. Main ideas
The strategy in this article follows from that in [chodosh-li-stryker]. A common feature in the proofs of Theorem 1.8 and Theorem 1.9 is the estimate on volume growth of ends by using $\mu$-bubbles, a technique introduced by Gromov [gromovmububble]. The method in [chodosh-li-stryker] cannot extend to $5$-dimensions directly because one cannot control curvature terms from the second fundamental form, the ambient Ricci curvature and the positive Jacobi function in the second variation of the warped $\mu$-bubble constructed there. The novelty of our method is to introduce a parameter (the exponent of the warped function) which brings more freedom to exploit curvature conditions. The same strategy works well in the resolution of stable Bernstein problem in $\mathbb{R}^{6}$ [mazet].
@@
Let $M^{n}\hookrightarrow(X^{n+1},g)$ $(4\leq n\leq 5)$ be a complete two-sided stable minimal hypersurface.
@@
    To show the vanishing of $A_{M}$ and $\operatorname{Ric}^{X}(\eta,\eta)$, we mainly prove that the hypersurface is parabolic. Then the standard equation $\Delta u+(|A_{M}|^{2}+\operatorname{Ric}^{X}(\eta,\eta))u=0$ will imply the result.
@@
    In dimension $6$, the fact that $M$ has at most one nonparabolic end follows from the Schoen–Yau inequality which requires only $\operatorname{biRic}_{X}\geq 0$ together with the infinite volume of $M$ , ensured by the bounded geometry assumption. When $\operatorname{Ric}^{X}_{3}\geq 0$, bounded geometry condition can be weaken to weakly bounded geometry; see [chodosh-li-stryker][Theorem 4.1] for details. In contrast, in dimension $5$, the splitting result in Theorem 3.1 implies that $M^{4}$ has at most two ends. Moreover, if $M^{4}$ has two ends, it must be a product manifold, see [antonelli-pozzetta-xu, catino-Mari-Mastrolia-Roncoroni, hong-wang-spectral-splitting]. Therefore, we only need to consider the case where $M^{4}$ has one end.
@@
    As mentioned in [chodosh-li-stryker], it is difficult to prove the linear volume growth of the nonparabolic end. Similarly, we consider an exhaustion $\{\Omega_{l}\}_{l=1}^{\infty}$ of $E$ and use the theory of $\mu$-bubbles to show $\partial\Omega_{l}\cap E$ has controlled volume. This is one difference comparing to [chodosh-li-stryker]. To control curvature terms in the second variation of the $\mu$-bubble, we introduce a parameter $k$ as the exponent of the warped function. Suitable choice of $k$ yields the desired volume estimate; see Lemma 4.7 in detail. This is the step that needs the assumption of strictly positive $k\operatorname{-triRic}$.
@@
    We need to upgrade volume control of the bubble to volume control of the slice which needs a lower bound on the intrinsic Ricci curvature of $M$. It is achieved by combining several curvature conditions and curvature estiamtes in Lemma 2.5.
@@
    Finally, to prove parabolicity, we need to construct test functions $\varphi_{i}$ such that $\int_{M}|\nabla\varphi_{i}|^{2}$ tends to zero. This is achievable due to the facts that we have at most one nonparabolic end and the nonparabolic end satisfies almost linear volume growth.
@@ 1.3. Organization
In Section 2, we recall some preliminaries on curvature estimates and ends of non-compact manifolds. In Section 3, we establish the splitting theorem and the one-end theorem under various curvature conditions. Section 4 is devoted to the balanced warped $\mu$-bubble exhaustion and volume growth estimates. The main results are proved in Section 5. The Appendix presents an example related to Theorem 1.9.
@@ 1.4. Acknowledgements
We would like to thank Prof. Xianzhe Dai, Prof. Martin Li, Prof. Guofang Wei and Prof. Rugang Ye for their interests in this work. The first author is supported by NSFC No. 12401058 and the Talent Fund of Beijing Jiaotong University No. 2024XKRC008. The second author is supported by a Simons Travel Grant.
@@ 2.1. Curvature conditions
If $(X^{n+1},g)$ is a Riemannian manifold, we denote the Riemann curvature tensor, the sectional curvature, the Ricci curvature tensor and the scalar curvature by $\operatorname{R}^{X}_{ijij}$, $\sec_{X}$, $\operatorname{Ric}^{X}_{ij}$ and $\operatorname{R}^{X}$ respectively. Let $M\hookrightarrow X^{n+1}$ be an immersion. We write $\operatorname{Ric}^{M}$ and $R^{M}$ to denote the Ricci curvature and scalar curvature of the pullback metric on $M$. We set the mean curvature to be the trace of the second fundamental form of the immersion.
@@
$|\sec_{X}|\leqslant\delta^{2},\quad\operatorname{inj}_{X}\geqslant\lambda$
@@
Let $(X^{n+1},g)$ be a Riemannian manifold and $\{e_{i}\}_{i=1}^{n+1}$ be an orthonormal basis for $T_{p}X$, $p\in X$. The bi-Ricci curvature along directions $(e_{1},e_{2})$ is defined by
$\operatorname{biRic}_{X}(e_{1},e_{2}):=\sum_{i=2}^{n+1}\operatorname{R}^{X}(e_{1},e_{i},e_{i},e_{1})+\sum_{j=3}^{n+1}\operatorname{R}^{X}(e_{2},e_{j},e_{j},e_{2}).$
We denote $\operatorname{biRic}_{X}:=\inf\{\operatorname{biRic}_{X}(e_{1},e_{2}):\ p\in X\ \text{and}\ e_{1},e_{2}\in T_{p}X\}.$
@@
Let $(X^{n+1},g)$ be a $n+1$-dimensional Riemannian manifold with bounded geometry. If $M^{n}\hookrightarrow(X^{n+1},g)$ is an immersed CMC hypersurface, then
$\operatorname{Ric}^{M}\geqslant-Cg,$
where the positive constant $C$ depends on $n$, $\sec_{X}$, $H^{M}$ and $\sup|A^{M}|$.
@@
Choose an orthonormal basis $\{e_{i}\}_{i=1}^{n+1}$ of $X^{n+1}$ at a point $p\in M$ with $e_{n+1}$ normal to $M$. The traced Gauss equation gives
$\operatorname{Ric}^{M}(e_{1},e_{1})=\sum_{i=1}^{n}\operatorname{R}^{X}_{i1i1}+A^{M}(e_{1},e_{1})H^{M}-\sum_{j=1}^{n}A^{M}(e_{1},e_{j})^{2},$
@@
Without bounded geometry, there is an alternative way to ensure the lower bound on $\operatorname{Ric}_{M}$.
@@
Let $(X^{n+1},g)$ be a $(n+1)$-dimensional Riemannian manifold with $\operatorname{Ric}^{X}_{n-1}\geq C_{1}$ for a constant $C_{1}$. If $M^{n}\hookrightarrow(X^{n+1},g)$ is an immersed CMC hypersurface, then
$\operatorname{Ric}^{M}\geqslant-Cg,$
where the positive constant $C$ depends on $n$, $C_{1}$, $H^{M}$ and $\sup|A^{M}|$. In particular, $C$ can be chosen to be zero if $M$ is totally geodesic and $C_{1}=0$.
@@
The result holds because
$\operatorname{Ric}^{M}(e_{1},e_{1})=$ $\sum_{i=1}^{n}\operatorname{R}^{X}_{i1i1}+A^{M}(e_{1},e_{1})H^{M}-\sum_{j=1}^{n}A^{M}(e_{1},e_{j})^{2}$
$\geq$ $\frac{5-n}{4}H_{M}^{2}-|A^{M}|^{2}$
where the inequality follows from
$|A^{M}|^{2}+H^{M}A^{M}(e_{1},e_{1})-\sum_{i=1}^{n}A^{M}(e_{1},e_{j})^{2}\geq\frac{5-n}{4}H_{M}^{2}.$
2.2. Curvature estimates
In this subsection, we recall the standard curvature estimates for stable CMC hypersurfaces in non-compact manifolds.
Lemma 2.5.
Let $(X^{n+1},g)$ $(4\leq n\leq 5)$ be a $(n+1)$-dimensional complete Riemannian manifold with weakly bounded geometry. Then there is a constant $C$ depending on $X$ and $H^{M}$ such that every complete noncompact stable CMC hypersurface (two-sided if it is minimal) $M^{n}\hookrightarrow(X^{n+1},g)$ with the mean curvature $H^{M}$ and compact boundary (possibly empty) satisfies
$\sup_{q\in\Omega}|A^{M}(q)|\min\{1,d_{M}(q,\partial\Omega)\}\leqslant C$
for any compact subset $\Omega\subset M$.
Proof.
Observe that when scaling the metric by the norm of the second fundamental form that tends to infinity, the limit of sequence of stable CMC hypersurface is a stable minimal hypersurface. Therefore, with the stable Bernstein theorem in $\mathbb{R}^{5}$ and $\mathbb{R}^{6}$, the statement follows from [chodosh-li-stryker][Lemma 2.4] and [Hong24][Lemma 3.6] with minor modification; see also [Rosenberg-Souam-Toubiana]. ∎
2.3. Ends of manifold
We now recall some preliminary results on parabolicity and nonparabolicity of manifolds; see [Li-Geometric-analysis] for more details.
@@ 2.3. Ends of manifold
Let $M$ be a complete noncompact manifold, and let $K\subset M$ be a compact subset and let $E$ be an unbounded connected component of $M\setminus K$ with smooth boundary. Then $E$ is said to be nonparabolic with respect to $K$ if there exists a positive harmonic function $f$ satisfying
$f|_{\partial E}=1,\ \text{and}\ \inf_{E}f=0.$
Otherwise, $E$ is parabolic.
Definition 2.7.
Fix a point $x\in M,$ and fix $L>0.$ A collection of open sets $\{E_{k}\}$ is an end of $M$ adapted to $x$ and length scale $L$ if each set $E_{k}$ is an unbounded connected component of $M\setminus\bar{B}_{kL}(x)$ and satisfies $E_{k+1}\subset E_{k}.$
Without lost of generality, we usually ignore the fixed point $x$ if it is clear in the context.
Definition 2.8.
A boundary end $\{E_{k}\}_{k=1}$ with length scale $L>0$ is nonparabolic if for every $k$, $E_{k}$ is an unbounded component of $M\backslash B_{kL}$ and $E_{k}$ is nonparabolic.
3. Ends of hypersurfaces
3.1. Splitting theorem and finite ends
In this subsection, we show that in a manifold $(X^{n+1},g)$, the number of ends of a stable CMC hypersurface $M^{n}$ is at most two under certain conditions.
The stability of $M$ implies that for any $\varphi\in C_{c}^{1}(M)$,
$\int_{M}|\nabla\varphi|^{2}\geq\int_{M}(|A|^{2}+\operatorname{Ric}^{X}(\nu,\nu))\varphi^{2}.$
Let $\operatorname{Ric}^{M}(x)=\inf_{x\in M}\{\operatorname{Ric}^{M}(\tau,\tau):|\tau|=1\ \text{and}\ \tau\in T_{x}M\}$. Suppose $e_{1}$ is an unit tangent vector of $M$ at $x$ such that $\operatorname{Ric}^{M}(x)=\operatorname{Ric}^{M}(e_{1},e_{1})$. Then the curvature rearrangement gives
$\int_{M}|\nabla\varphi|^{2}+\alpha\operatorname{Ric}^{M}(x)\varphi^{2}\geq$ $\int_{M}(|A|^{2}+\operatorname{Ric}^{X}(\nu,\nu)+\alpha\operatorname{Ric}^{M}(e_{1},e_{1}))\varphi^{2}$
$\geq$ $\int_{M}(\operatorname{Ric}^{X}(\nu,\nu)+\alpha\sum_{i=1}^{n}\operatorname{R}^{X}_{1i1i})\varphi^{2}$
$+$ $\int_{M}(|A|^{2}+\alpha A_{11}H-\alpha\sum_{i=1}^{n}A^{2}_{1i})\varphi^{2}.$
Terms involving the second fundamental from on the right-hand side is bounded from below by
$\frac{4-(n-1)\alpha^{2}}{4((1-\alpha)n+\alpha)}H^{2},\quad\alpha<\frac{n}{n-1};$
see Deng [dengqintao, Lemma 2.1]. Moreover, the equality holds iff
$A_{11}=-\frac{(n-1)\alpha-2}{2((1-\alpha)n+\alpha)}H,\quad A_{ii}=\frac{2-\alpha}{2((1-\alpha)n+\alpha)}H,\quad i\neq 1.$
Assuming that $\alpha\text{-}\operatorname{biRic}_{X}(e_{1},\nu):=\operatorname{Ric}^{X}(\nu,\nu)+\alpha\sum_{i=1}^{n}\operatorname{R}^{X}_{1i1i}$ satisfies
$\alpha\text{-}\operatorname{biRic}+\frac{4-(n-1)\alpha^{2}}{4((1-\alpha)n+\alpha)}H^{2}\geq 0,$
we have
$\int_{M}|\nabla\varphi|^{2}+\alpha\operatorname{Ric}^{M}(x)\varphi^{2}\geq 0.$
Then it follows from works of Antonelli–Pozzetta–Xu [antonelli-pozzetta-xu], Catino–Mari–Mastrolia–Roncoroni [catino-Mari-Mastrolia-Roncoroni] and Hong–Wang [hong-wang-spectral-splitting] (using different ideas) that if $M$ has at least two ends and $\alpha>\frac{n-1}{4}$, then $\operatorname{Ric}^{M}\geq 0$ on $M.$ By the classical Cheeger–Gromoll splitting theorem, $M$ is a product manifold $\Sigma\times\mathbb{R}$ with linear volume growth, and thus parabolic manifold.
Thus, we have the following theorem.